Find the differential equation of the space curve in which the two families of surfaces

u = x2 − y2 = c1 and

v = y2 − z2 = c2 intersect.


1
Expert's answer
2021-03-09T05:34:22-0500

"du=2xdx-2ydy=0"

"xdx=ydy"

"dv=2ydy-2zdz=0"

"ydy=zdz"

"xdx=ydy=zdz"

"\frac{xdx}{xyz}=\frac{ydy}{xyz}=\frac{zdz}{xyz}"

"\frac{dx}{yz}=\frac{dy}{xz}=\frac{dz}{xy}"

"yzp+xzq=xy"